OpenStudy (anonymous):

A tent has a base that is 20ft. across and a slant height of 14ft. To the nearest whole unit, what is the area of the tent , including the floor. use 3.14 for p

OpenStudy (s):

Surface area of a cone = pi * r * s + pi * r^2 pi = 3.14 r = d/2 = 20/2 = 10 s = 14 SA = 3.14 * 10 * 14 + 3.14 * 10^2 SA = 439.6 + 314 SA = 753.6 = 754 sq ft